OpenStudy (anonymous):

need to integrate: (4x^(1/2))/(x+16)

OpenStudy (kainui):

Is this your integral? \[\LARGE \int\limits \frac{4 \sqrt{x}}{x+16}dx\]

OpenStudy (anonymous):

yes

OpenStudy (kainui):

It looks like trig substitution will work here if you pick \[\LARGE x=16 \tan^2 \theta\]